# Exploring the hardness of the ionizing radiation with the infrared softness diagram

### v1

@@ 5.3 An interpretation of the bimodal distributions
The results described above clearly indicate the existence of two well-defined families of AGN in the Universe, based on the hardness of their radiation field and the excitation of the gas in the NLR surrounding the active nuclei. These two families emerge even when a single softness parameter is analysed, thanks to the high statistical significance of the IDEOS sample and the use of emission lines covered by Spitzer. Moreover, the families are better characterized using the HCm-Teff-IR code, which derives parameters uniquely tracing the hardness of the radiation field, such as , while simultaneously providing parameters that trace gas excitation, such as log .
@@
In the framework of AGN classification, the spectral types are commonly divided into type 1 and type 2 nuclei, as described by the unified model. Type 1 AGN are characterized by broad emission lines (e.g., H and Mg ii) in their optical and ultraviolet spectra, which arise from high-velocity gas in the broad-line region (BLR) near the central supermassive black hole. In contrast, type 2 AGN exhibit only narrow emission lines, as the BLR is obscured by dust when viewed from certain angles. This geometric interpretation has been a cornerstone of AGN taxonomy, providing a unifying framework to explain the different observed properties of these sources. However, a geometric interpretation of the spectral properties does not fully account for the physical differences in radiation field hardness and gas excitation.
@@
Our findings challenge the traditional classification schemes, suggesting the presence of two distinct physical states in AGN, beyond the type 1/2 dichotomy. One family exhibits hard radiation fields combined with low-excitation gas (HRLE), while the other shows soft radiation fields and highly excited gas (SRHE). These two families are robustly identified and do not depend on the assumed physical conditions in the models. By examining mid- and far-IR emission lines, we aim to elucidate the underlying mechanisms responsible for the existence of these two AGN families.
@@
On the one hand, one may question to what extent these results can be affected by the contribution of possible star-formation emission, a critical point given the low spatial resolution provided by Spitzer. Taking advantage of the extensive database of spectral measurements available in the IDEOS catalog, we investigated a possible dependence of or log on the contribution from star formation activity. For this purpose we used the equivalent widths of certain PAH bands observed in the mid-IR, such as those at 6.2 m and 11.3 m. Although the strong radiation field in the vicinity of AGN is expected to destroy the complex molecules responsible for these emission bands (e.g., Smith et al. 2007; Diamond-Stanic & Rieke 2010), a residual contribution could still be present. Thus, we explored possible correlations in the studied sample between the observed EWs of the PAH bands and the parameters derived by our code HCm-Teff-IR, shown in Fig. 6, assuming models with dust grains and a stopping criterion of . In any case, the results obtained under different conditions are practically the same, meaning that the assumptions made in the models do not alter our results. No clear correlation is obtained for any of the four relations. A very slight, but not significant, correlation coefficient () is found for the relation between and the EW of the two bands (around -0.20), indicating a lower EW value for higher , but this is far from conclusive. Furthermore, this correlation is mainly observed for objects belonging to the family with a lower (i.e., around -1.7), while no trace of correlation is detected for objects with higher . Additionally, no correlation is observed for , as shown in Fig. 6. A similar result is obtained when the EWs are represented against the parameters. A large dispersion is also observed in subsamples grouped by AGN type, which show no apparent correlation.
@@
Given the lack of correlation between the hardness of the radiation field in AGN and the presence of PAHs, we conclude that or log are not affected by the contribution of star formation in our sample. Additionally, it is reasonable to consider that the spatial resolution of Spitzer/IRS observations may be insufficient, for most galaxies, to resolve the regions where PAHs are destroyed by AGN radiation from those where UV photons escaping from central galactic zones excite the vibrational modes of these molecules, thereby increasing the intensity of the PAH bands.
@@
Therefore, the causes behind the observed bimodal distribution in the AGN radiation field should be sought in the AGN’s own properties. In this way, our result can be interpreted as an evidence that the slope of the ionizing continuum in AGN – and therefore the SED shape – does not exhibit a flat or smooth transition from soft to hard continua. Instead, it shows a dual-mode behavior, where active nuclei have either low- or high-hardness properties. This dichotomy was noted in previous studies of radio galaxy samples (Laing et al., 1994; Buttiglione et al., 2010). However, our analysis extends this finding to the broader AGN population. The bimodal behavior found in this study aligns with the two main accretion states identified in AGN (Körding et al., 2006; Svoboda et al., 2017; Fernández-Ontiveros & Muñoz-Darias, 2021), suggesting that the ionizing continuum in high-excitation sources is powered by the accretion disk, which dominates the energy output in accreting black holes during the so-called soft state or radiatively efficient mode. Conversely, low-excitation sources would be associated with the hard state or kinetic mode, characterized by a cooler and/or truncated disk, which may lead to a non-thermal continuum dominated by the jet or corona (e.g. Capetti et al., 2005; Fernández-Ontiveros et al., 2023). The observed bimodality in both and could then be attributed to the rapid transition between these two main accretion states.
In addition to the hardness bimodality, it is worth exploring the reasons behind the simultaneous bimodality in the NLR gas excitation, and, most importantly, why the latter shows strong statistical significance in its correlation with the radiation field hardness. Specifically, most HR objects exhibit LE, while most SR objects have HE.
@@
To further investigate this effect, we examined the spectral classification of the galaxies in our sample. Although these are not available for all galaxies, they may reveal certain trends. Table LABEL:class provides the number of galaxies assigned to different spectral main IR categories as defined by Spoon et al. (2022): 1A have hot-dust-dominated spectra typical of AGN; 1B have IR spectra similar to those found in optical Seyferts; and 2A-3B are ULIRG, mainly in interacting systems. Additionally, the table shows the galaxy distribution across the four resulting groups formed by the combinations of the two marginal distributions for and log (i.e. SRLE, SRHE, HRLE, and HRHE). These classifications were based on models with = 99.9% including dust grains, though the results are not significantly different when other model conditions are assumed. From this classification, it seems that no clear trend is found for IR-bright galaxies, such as 1A or 2A-3B types, which are more evenly distributed between the two main categories. On the other hand, it could seem that 1B galaxies, that includes Seyfert as identified from their IR spectra, are mainly classified as SRHE. Table LABEL:class also includes the median luminosity for each AGN type subsample shown in Fig. 5. Seyfert 1 and 1B nuclei show the brightest luminosities (), although the other spectral types show similar values , except for LINERs, which are clearly fainter sources ().
@@
This trend is partially confirmed with the results coming from the independent optical classification, also shown in Table LABEL:class and available in the objects compiled by Fernández-Ontiveros et al. (2016), and complemented for a substantial number of the objects of the IDEOS catalog in Véron-Cetty & Véron (2010). This classification points to a larger number of SRHE in the Sy1 type, while intermediate types (S1.x) are more equally distributed between SRHE and HRLE. On the other hand, a larger number of HRLE are found for Sy2, hidden broad-line Seyferts (Sy1h) and LINERs. Nevertheless, a non-negligible number of LINERs lie in one of the intermediate categories, such as SRLE, which could indicate that they can be more easily found in a transition process. However, the scarce number of objects classified in this category prevents us from extracting more robust conclusions. Thus, despite the apparently contradictory results for Sy2 from the optical and IR classifications, the statistical trend to find with a larger probability AGN with prominent broad emission lines in the SRHE phase underlines the existence of a physical basis related with the variations of the accretion rate in the innermost disk, though, purely phenomenological classifications based solely on certain optical or IR spectral features are not necessarily representative of the accretion state of the active nucleus.
@@
Moreover, the trend to identify LINERs as HRLE is consistent with several studies (e.g. Ho 1999; Maoz 2007; Vagnetti et al. 2013) based on direct X-ray observations, which suggest that LINERs have a weaker blue bump luminosity combined with a less efficient radiative rate in their accretion disks (e.g. Gliozzi et al. 2008). This is consistent with the lower median bolometric luminosity found among the LINER subsample using the [Ne v] 14.3 m line (see Table LABEL:class). As a result, typical values of for LINERs lie between -1.2 and -1.6, while values around -1.5 and -1.8 have been reported for Seyfert galaxies (e.g. Constantin et al. 2009; Eracleous et al. 2010; Younes et al. 2011; Hernández-García et al. 2013), in agreement with our results obtained from the IR classification and for a substantial fraction of the sample as compared with the optical classification.
@@
The different causes proposed to explain the lower gas excitation observed in LINERs include the presence of high-velocity shocks, that can enhance low-excitation emission lines (e.g. Dopita et al. 1997; Shields & Sabra 2003), a possible higher gas density, which could imply a higher fraction of particles relative to the number of ionizing photons (e.g. Filippenko & Halpern 1984), or gas geometry, as a possible ionizing photon escape fraction could affect the balance between photons and particles (e.g. Binette et al. 1996; Stasińska et al. 2008). Nonetheless, it is difficult to establish a clear relation between the first two causes and the observed bimodality found in the analysis of mid- and far-IR lines. In contrast, the third scenario appears more plausible, considering that a sudden hardening of the radiation field within a limited-volume gas distribution can convert a radiation-bounded geometry, in which almost all photons are absorbed, into a matter-bounded one, where the most energetic photons pass through without being absorbed due to the much lower cross section of ions for high-energy photons. Thus, a change in the nature of the accretion disk could also result in a different ionization structure of the surrounding gas, leading to a much lower value for as increases.
In any case, more data is required but, thanks to the new ongoing and future IR facilities for low redshift, such as JWST or METIS, and for high redshift, such as ALMA, this bimodality can be explored in greater detail to determine whether a population exists that fills the gap in the bimodality observed at low redshift, or if the rapid transition is confirmed.
@@ 6 Summary and conclusions
In this paper, we analyzed a sample of AGN-selected galaxies with available emission-line information in the mid- and far-IR rest-frame spectral regimes, in order to use adequate emission-line ratios involving high-excitation lines, such as [Ne v] or [O iv], to trace the shape of the hardness of the incident radiation field. We defined two forms of this parameter, using the ratio of Ne23Ne5 to O3O4, and adapting it for a broader use by including a parameter exclusively based on mid-IR lines, S34O4. Compared to large grids of photoionization models, the newly defined parameters are sensitive to both the , which traces the slope of the AGN continuum between the X-ray and UV ranges, and log , an indicator of the gas excitation. The study of these new softness parameters designed for AGN reveals the existence of a clear bimodal distribution in our selected sample. This bimodal behavior is further confirmed by the application of the HCm-Teff-IR code, which provides the corresponding distributions of and log .
Moreover, the existence of two well-defined AGN families is established, concentrated around either a soft radiation field (SR with -1.7) and highly excited gas in the NLR (HE with log -1.5), or a hard radiation field (HR with -1.4) and low-excitation gas (LE with log -2.4). These two categories are identified regardless of the geometry and presence of dust grains assumed in the model grids used for comparison with the observed emission-line ratios.
The existence of a bimodality in both the hardness of the radiation field as seen in the NLR of AGN and derived from mid- and far-IR lines, and in the ionization parameter, is consistent with a dual scenario of accretion efficiency in black holes, along with variable geometry of the surrounding gas as a function of the spectral energy distribution emitted by the disk.
@@ Acknowledgements.
We thank the anonymous referee, whose very helpful and constructive comments and suggestions have helped to improve our manuscript. This work has been funded by project Estallidos8 PID2022-136598NB-C32 (Spanish Ministerio de Ciencia e Innovación). We also acknowledge financial support from the State Agency for Research of the Spanish MCIU through the ”Center of Excellence Severo Ochoa” award to the Instituto de Astrofísica de Andalucía (SEV-2017-0709). JAFO acknowledges financial support by the Spanish Ministry of Science and Innovation (MCIN/AEI/10.13039/501100011033), by “ERDF A way of making Europe” and by “European Union NextGenerationEU/PRTR” through the grants PID2021-124918NB-C44 and CNS2023-145339; MCIN and the European Union – NextGenerationEU through the Recovery and Resilience Facility project ICTS-MRR-2021-03-CEFCA. EPM also acknowledges the assistance from his guide dog Rocko without whose daily help this work would have been much more difficult.References

### v2

@@ 5.3 An interpretation of the bimodal distributions
The results described above clearly indicate the existence of two well-defined families of AGNs in the Universe, based on the hardness of their radiation field and the excitation of the gas in the NLR surrounding the active nuclei. These two families emerge even when a single softness parameter is analyzed, thanks to the high statistical significance of the IDEOS sample and the use of emission lines covered by Spitzer. Moreover, the families are better characterized using the HCm-Teff-IR code, which derives parameters uniquely tracing the hardness of the radiation field, such as , while simultaneously providing parameters that trace gas excitation, such as log .
@@
In the framework of AGN classification, the spectral types are commonly divided into type 1 and type 2 nuclei, as described by the unified model. Type 1 AGNs are characterized by broad emission lines (e.g., H and Mg ii) in their optical and ultraviolet spectra, which arise from high-velocity gas in the broad-line region (BLR) near the central supermassive black hole. In contrast, type 2 AGNs exhibit only narrow emission lines, as the BLR is obscured by dust when viewed from certain angles. This geometric interpretation has been a cornerstone of AGN taxonomy, providing a unifying framework to explain the different observed properties of these sources. However, a geometric interpretation of the spectral properties does not fully account for the physical differences in radiation field hardness and gas excitation.
@@
Our findings challenge the traditional classification schemes, suggesting the presence of two distinct physical states in AGNs, beyond the type 1/2 dichotomy. One family exhibits hard radiation fields combined with low-excitation gas (HRLE), while the other shows soft radiation fields and highly excited gas (SRHE). These two families are robustly identified and do not depend on the assumed physical conditions in the models. By examining mid- and far-IR emission lines, we aim to elucidate the underlying mechanisms responsible for the existence of these two AGN families.
@@
On the one hand, one may question to what extent these results can be affected by the contribution of possible star-formation emission, a critical point given the low spatial resolution provided by Spitzer. Taking advantage of the extensive database of spectral measurements available in the IDEOS catalog, we investigated a possible dependence of or log on the contribution from star formation activity. For this purpose we used the equivalent widths of certain PAH bands observed in the mid-IR, such as those at 6.2 m and 11.3 m. Although the strong radiation field in the vicinity of AGNs is expected to destroy the complex molecules responsible for these emission bands (e.g., Smith et al. 2007; Diamond-Stanic & Rieke 2010), a residual contribution could still be present. Thus, we explored possible correlations in the studied sample between the observed EWs of the PAH bands and the parameters derived by our code HCm-Teff-IR, shown in Fig. 6, assuming models with dust grains and a stopping criterion of . In any case, the results obtained under different conditions are practically the same, meaning that the assumptions made in the models do not alter our results. No clear correlation is obtained for any of the four relations. A very slight, but not significant, correlation coefficient () is found for the relation between and the EW of the two bands (around -0.20), indicating a lower EW value for higher , but this is far from conclusive. Furthermore, this correlation is mainly observed for objects belonging to the family with a lower (i.e., around -1.7), while no trace of correlation is detected for objects with higher . Additionally, no correlation is observed for , as shown in Fig. 6. A similar result is obtained when the EWs are represented against the parameters. A large dispersion is also observed in subsamples grouped by AGN type, which show no apparent correlation.
@@
Given the lack of correlation between the hardness of the radiation field in AGNs and the presence of PAHs, we conclude that or log are not affected by the contribution of star formation in our sample. Additionally, it is reasonable to consider that the spatial resolution of Spitzer/IRS observations may be insufficient, for most galaxies, to resolve the regions where PAHs are destroyed by AGN radiation from those where UV photons escaping from central galactic zones excite the vibrational modes of these molecules, thereby increasing the intensity of the PAH bands.
@@
Therefore, the causes behind the observed bimodal distribution in the AGN radiation field should be sought in the AGN’s own properties. In this way, our result can be interpreted as evidence that the slope of the ionizing continuum in AGNs – and therefore the SED shape – does not exhibit a flat or smooth transition from soft to hard continua. Instead, it shows a dual-mode behavior, where active nuclei have either low- or high-hardness properties. This dichotomy was noted in previous studies of radio galaxy samples (Laing et al., 1994; Buttiglione et al., 2010). However, our analysis extends this finding to the broader AGN population. The bimodal behavior found in this study aligns with the two main accretion states identified in AGNs (Körding et al., 2006; Svoboda et al., 2017; Fernández-Ontiveros & Muñoz-Darias, 2021), suggesting that the ionizing continuum in high-excitation sources is powered by the accretion disk, which dominates the energy output in accreting black holes during the so-called soft state or radiatively efficient mode. Conversely, low-excitation sources would be associated with the hard state or kinetic mode, characterized by a cooler and/or truncated disk, which may lead to a non-thermal continuum dominated by the jet or corona (e.g., Capetti et al., 2005; Fernández-Ontiveros et al., 2023). The observed bimodality in both and could then be attributed to the rapid transition between these two main accretion states.
In addition to the hardness bimodality, it is worth exploring the reasons behind the simultaneous bimodality in the NLR gas excitation and, most importantly, why the latter shows strong statistical significance in its correlation with the radiation field hardness. Specifically, most HR objects exhibit LE, while most SR objects have HE.
@@
To further investigate this effect, we examined the spectral classification of the galaxies in our sample. Although these are not available for all galaxies, they may reveal certain trends. Table LABEL:class provides the number of galaxies assigned to different spectral main IR categories as defined by Spoon et al. (2022): 1A have hot-dust-dominated spectra typical of AGNs; 1B have IR spectra similar to those found in optical Seyferts; and 2A-3B are ULIRG, mainly in interacting systems. Additionally, the table shows the galaxy distribution across the four resulting groups formed by the combinations of the two marginal distributions for and log (i.e., SRLE, SRHE, HRLE, and HRHE). These classifications were based on models with = 99.9% including dust grains, though the results are not significantly different when other model conditions are assumed. From this classification, it seems that no clear trend is found for IR-bright galaxies, such as 1A or 2A-3B types, which are more evenly distributed between the two main categories. On the other hand, it could seem that 1B galaxies, that includes Seyfert as identified from their IR spectra, are mainly classified as SRHE. Table LABEL:class also includes the median luminosity for each AGN type subsample shown in Fig. 5. Seyfert 1 and 1B nuclei show the brightest luminosities (), although the other spectral types show similar values , except for LINERs, which are clearly fainter sources ().
@@
This trend is partially confirmed with the results coming from the independent optical classification, also shown in Table LABEL:class and available in the objects compiled by Fernández-Ontiveros et al. (2016), and complemented for a substantial number of the objects of the IDEOS catalog in Véron-Cetty & Véron (2010). This classification points to a larger number of SRHE in the Sy1 type, while intermediate types (S1.x) are more equally distributed between SRHE and HRLE. On the other hand, a larger number of HRLE are found for Sy2, hidden broad-line Seyferts (Sy1h) and LINERs. Nevertheless, a non-negligible number of LINERs lie in one of the intermediate categories, such as SRLE, which could indicate that they can be more easily found in a transition process. However, the scarce number of objects classified in this category prevents us from extracting more robust conclusions. Thus, despite the apparently contradictory results for Sy2 from the optical and IR classifications, the statistical trend to find with a larger probability AGNs with prominent broad emission lines in the SRHE phase underlines the existence of a physical basis related with the variations of the accretion rate in the innermost disk, though, purely phenomenological classifications based solely on certain optical or IR spectral features are not necessarily representative of the accretion state of the active nucleus.
@@
Moreover, the trend to identify LINERs as HRLE is consistent with several studies (e.g., Ho 1999; Maoz 2007; Vagnetti et al. 2013) based on direct X-ray observations, which suggest that LINERs have a weaker blue bump luminosity combined with a less efficient radiative rate in their accretion disks (e.g., Gliozzi et al. 2008). This is consistent with the lower median bolometric luminosity found among the LINER subsample using the [Ne v] 14.3 m line (see Table LABEL:class). As a result, typical values of for LINERs lie between -1.2 and -1.6, while values around -1.5 and -1.8 have been reported for Seyfert galaxies (e.g., Constantin et al. 2009; Eracleous et al. 2010; Younes et al. 2011; Hernández-García et al. 2013), in agreement with our results obtained from the IR classification and for a substantial fraction of the sample as compared with the optical classification.
@@
The different causes proposed to explain the lower gas excitation observed in LINERs include the presence of high-velocity shocks, that can enhance low-excitation emission lines (e.g., Dopita et al. 1997; Shields & Sabra 2003), a possible higher gas density, which could imply a higher fraction of particles relative to the number of ionizing photons (e.g., Filippenko & Halpern 1984), or gas geometry, as a possible ionizing photon escape fraction could affect the balance between photons and particles (e.g., Binette et al. 1996; Stasińska et al. 2008). Nonetheless, it is difficult to establish a clear relation between the first two causes and the observed bimodality found in the analysis of mid- and far-IR lines. In contrast, the third scenario appears more plausible, considering that a sudden hardening of the radiation field within a limited-volume gas distribution can convert a radiation-bounded geometry, in which almost all photons are absorbed, into a matter-bounded one, where the most energetic photons pass through without being absorbed due to the much lower cross section of ions for high-energy photons. Thus, a change in the nature of the accretion disk could also result in a different ionization structure of the surrounding gas, leading to a much lower value for as increases. In any case, more data are required, but thanks to the new ongoing and future IR facilities for low redshift (such as JWST or METIS) and for high redshift (such as ALMA), this bimodality can be explored in greater detail to determine whether a population exists that fills the gap in the bimodality observed at low redshift or if the rapid transition is confirmed.
@@ 6 Summary and conclusions
In this paper, we have analyzed a sample of AGN-selected galaxies with available emission-line information in the mid- and far-IR rest-frame spectral regimes in order to use adequate emission-line ratios involving high-excitation lines, such as [Ne v] or [O iv], to trace the shape of the hardness of the incident radiation field. We defined two forms of this parameter using the ratio of Ne23Ne5 to O3O4 and adapting it for a broader use by including a parameter exclusively based on mid-IR lines, S34O4. Compared to large grids of photoionization models, the newly defined parameters are sensitive to both the , which traces the slope of the AGN continuum between the X-ray and UV ranges, and log , an indicator of the gas excitation. Our study of these new softness parameters designed for AGNs revealed the existence of a clear bimodal distribution in our selected sample. This bimodal behavior is further confirmed by the application of the HCm-Teff-IR code, which provides the corresponding distributions of and log .
Moreover, the existence of two well-defined AGN families has been established, which are concentrated around either a SR continuum (with -1.7) and HE gas in the NLR (with log -1.5) or a HR continuum (with -1.4) and LE gas (with log -2.4). The presence of these two categories can be identified regardless of the geometry and presence of dust grains assumed in the model grids used for comparison with the observed emission-line ratios. The existence of a bimodality in both the hardness of the radiation field, as seen in the NLR of AGNs and derived from mid- and far-IR lines, and in the ionization parameter is consistent with a dual scenario of accretion efficiency in black holes and variable geometry of the surrounding gas as a function of the SED emitted by the disk.
@@ Acknowledgements.
We thank the anonymous referee, whose very helpful and constructive comments and suggestions have helped to improve our manuscript. This work has been funded by project Estallidos8 PID2022-136598NB-C32 (Spanish Ministerio de Ciencia e Innovación). We also acknowledge financial support from the State Agency for Research of the Spanish MCIU through the “Center of Excellence Severo Ochoa” award to the Instituto de Astrofísica de Andalucía (SEV-2017-0709). JAFO acknowledges financial support by the Spanish Ministry of Science and Innovation (MCIN/AEI/10.13039/501100011033), by “ERDF A way of making Europe” and by “European Union NextGenerationEU/PRTR” through the grants PID2021-124918NB-C44 and CNS2023-145339; MCIN and the European Union – NextGenerationEU through the Recovery and Resilience Facility project ICTS-MRR-2021-03-CEFCA. EPM also acknowledges the assistance from his guide dog Rocko without whose daily help this work would have been much more difficult.References